A car accelerates from rest at a constant rate α for some time after which it decelerates at a constant rateβ and comes to rest. if total time elasped is t, then maximum velocity acquired by car will be
1
Expert's answer
2019-07-01T09:28:42-0400

The maximum velocity acquired by car:


"v=\alpha t_1=\beta t_2"

"t=t_1+ t_2"

Thus,


"t=\frac{v}{\alpha}+\frac{v}{\beta}"

"v=\frac{t}{\frac{1}{\alpha}+\frac{1}{\beta}}"